# L3Ms — Lagrange Large Language Models

### v1

@@ 5 Lagrange Large Language Models (L3Ms)
We have thus far formulated the SFT and alignment stages as a constrained optimization problem in Eq. 5. We proceed to find solutions for the same by solving the unconstrained objective in Eq. 10. We call the family of models obtained in this way L3Ms, i.e., Lagrange Large Language Models.
@@ 5.1 Optimization procedure
Since the log-barrier converges to the characteristic function as $\mu\rightarrow 0$, we want to find the minimizer of $\mathcal{G}_{\mu}(\theta)$ for a very small $\mu$. However, doing so directly leads to instabilities as the objective function is ill-conditioned. Instead, it is common practice to follow an iterative procedure where one finds the minimizer for a fixed $\mu$, reduces $\mu$, and repeats (Curtis et al., 2024). Specifically, the procedure is instantiated with initial values $\theta_{0}$, $\mu_{1}$, and $10<\gamma<$ 1. On the $t$-th iteration, $\mu_{t}\leftarrow\gamma\mu_{t-1}$ is reduced and $\theta_{t}\leftarrow\arg_{\theta}\min\mathcal{G}_{\mu_{t}}(\theta)$ (with the initialization at $\theta_{t-1}$). In doing so, the constraints are enforced gradually, nudging the LLM to satisfy them over the training procedure while avoiding instabilities. Furthermore, as $\{\mu_{t}\}\searrow 0$, the weights $\theta_{t}$ converges to the minimizer of the constrained problem.
It is impossible to minimize $\mathcal{G}_{\mu_{t}}(\theta)$ exactly in many practical applications. Instead, at each iteration, we take a single optimization step toward the solution. Doing so is amenable to stochastic gradient methods and mitigates computational overhead — the training proceeds as normal while the value of $\mu$ is reduced over the course of the training. One can guarantee convergence of such a procedure to the optimal solution in some settings; for instance, Curtis et al. (2024) prove convergence when dealing with box constraints. However, convergence in a scenario like ours is not guaranteed. Nonetheless, we will experimentally demonstrate its use for our constrained problems.
@@
This follows immediately from Eqs. 9 and 10. Note that the $\partial_{\theta}C_{i}(\theta)$’s are also known as the policy gradients in reinforcement learning literature. We discuss our strategy for estimating these gradients in Appendix B and refer readers to Schulman et al. (2016) for a more detailed review.
@@ 5.2 Connection to Lagrange multipliers
The log-barrier and the Lagrangian are intrinsically connected; this becomes evident when comparing Eq. 11 with the (gradient of the) Lagrangian in Eq. 8. In particular, we define the multipliers:
@@ 5.3 Implementation details
Alternating objective and gradient clipping.
A simple, yet effective way to ensure stable training of large models is gradient clipping (Goodfellow et al., 2016; Zhang et al., 2020). We employ this technique, albeit with the modification of clipping the gradients of both the task objective and the constraints separately, as they can have varying magnitudes. We achieve this by alternating between reducing the task objective and enforcing the constraints by flipping a fair coin to select one or the other. While this doubles the number of steps to achieve the same effect, it does not increase the amount of work done as now only one part of the objective or the other is evaluated at each step.
Length normalization.
The gradient of our objective function in Eq. 11 involves the log-likelihoods of the generated responses through the gradients $\partial_{\theta}C_{i}(\theta)$’s (cf. Eq. 12). To avoid a response length bias, we length-normalize the log-likelihoods, akin to the definition of perplexity.
Estimating the mean preference rewards.
We need to estimate the expectations involved in the gradient of our objective function in Eq. 11. The expectations in the numerators can be estimated with the per-mini-batch Monte Carlo averages (Mohamed et al., 2020). However, $C_{i}(\theta)$ in the denominator needs careful consideration. Note that: (i) $C_{i}(\theta)$ does not involve the gradient, so its estimates can include information from previous mini-batches to reduce variance, and (ii) since the weights are updated during training, $C_{i}(\theta)$ is non-stationary. Consequently, we use an exponential moving average estimate for the mean (offset) preference rewards $C_{i}(\theta)$ in the denominator.
@@ 6 Experimental results
To illustrate the versatility of L3Ms, we empirically evaluate them on two criteria: (i) the satisfaction of the imposed custom constraints, and (ii) the task perplexity objective. Note that L3Ms have essentially the same time complexity as traditional approaches: the SFT and alignment stages have been combined into one and the log-barrier parameter is adjusted during the training procedure itself.
@@ 6.1 Setup
We use LLaMA-7B (Touvron et al., 2023) for all our experiments, as it is a lightweight LLM pre-trained on a large corpus. We are interested in the instruction-following task, for which we use UltraChat (Ding et al., 2023), a large-scale dataset of instructional conversations. We refer to the model trained via the SFT objective (without alignment) as the SFT model. We train LLMs using the min-max approach to find the saddle-point of the Lagrangian, as proposed by Moskovitz et al. (2024); Dai et al. (2024), referring to them as MMs. Lastly, we refer to models trained using our approach as L3Ms. We use the Transformers package (Wolf et al., 2020) for our implementation and run experiments on NVIDIA H100s. Further details on our setup are included in Appendix A.
@@ 6.2 Length constrained L3Ms
Consider tasks where the lengths of the responses need to be contained in the range $[l_{\text{low}},l_{\text{high}}]$ to control its verbosity, in summarization tasks for example (Makino et al., 2019). The natural choice of reward functions in this case are the ones that compute the response length and its negation via $r_{1}(y|x)=|y|$ and $r_{2}(y|x)=-|y|$; correspondingly, these rewards are to be controlled with the requirements of $l_{\text{low}}$ and $l_{\text{high}}$ respectively. Note that these rewards are perfectly anti-correlated.
If we naively average the rewards, any unconstrained formulation of alignment (including RLHF) will be ineffective as the loss will always vanish due to the anti-correlation. We could use a weighted average and tune the weights heuristically, but this is tedious. Instead, we use the constrained formulation and directly constrain the rewards $r_{1}(y|x)=|y|\geq l_{\text{low}}$ and $r_{2}(y|x)=-|y|\geq-l_{\text{high}}$.
We train several L3Ms with varying length constraints. We illustrate the distributions of the generated response lengths (in tokens) and report the perplexities achieved on the task-related data in Fig. 2. We observe that the mean response lengths are in the required range in each case, satisfying the imposed expectation constraints. Additionally, the task perplexities increase slightly as the constraints on the response lengths are made more stringent. However, there is little to no degradation relative to the SFT model, with all mean task perplexities being within 0.02 standard deviations.
The examples included in Table 1 are generated from such L3Ms. While all the responses correctly answer the prompts, their lengths vary corresponding to the length constraints imposed on them.
@@
Figure 2: Length constrained L3Ms. We report the response lengths (in tokens) and task perplexities of the SFT model and the L3Ms with varying length constraints. Left: The mean response length with the mean and standard deviation of the task perplexities. Right: The distribution of the response lengths. The notches indicate the medians and their 95% confidence intervals, the boxes show the $\pm$25quantiles, and the whiskers denote the 1.5$\times$ interquartile ranges. The white circles mark the means and the black dashed lines depict the constraints imposed on the different L3Ms.
@@ 6.3 Helpful and harmless L3Ms
Next, we consider the Helpful and Harmless preferences that have been extensively used in the LLM alignment literature (Ji et al., 2023; Wang et al., 2024; Zhou et al., 2024; Guo et al., 2024). Specifically, we utilize the datasets by Bai et al. (2022) to train two preference reward functions respectively. These learned reward functions are negatively correlated (Bai et al., 2022; Dai et al., 2024). Furthermore, note that the numerical outputs of both these reward functions are interpreted as ratings such that a higher numerical value indicates higher helpfulness/harmlessness and vice versa.
We train several L3Ms with varying HH constraints. We compare our L3M approach of using log-barriers with the min-max optimization by Moskovitz et al. (2024); Dai et al. (2024) to find the saddle-point of the Lagrangian (MMs). In our experience, learning the Lagrange multipliers in this way is extremely sensitive to the choice of the learning rate. Moreover, to avoid loading an additional reference LLM during the training process, as is done by Moskovitz et al. (2024); Dai et al. (2024), our implementation of MM minimizes the task perplexity directly (as is done by L3Ms as well).
Fig. 3 shows the achieved task perplexities and helpful-harmless rewards for the different LLMs. At initialization, the helpful and harmless rewards are both low, with a high task perplexity of 1.316$\pm$0.4. This is improved upon by the SFT model, reducing the perplexity to 0.805$\pm$0.3 and attaining a high helpfulness reward (attributed to the task data instilling instruction-following capabilities). Further, MMs sacrifice task performance to over-satisfy the constraints, with mean task perplexities higher than 0.822. On the other hand, L3Ms satisfy the imposed helpful-harmless reward constraints with lower task perplexities (the mean perplexities are in the range of 0.804-0.820). We attribute this to the L3Ms having better Lagrange multipliers by design rather than learning them as is done in MMs.
@@ 7 Conclusions
In this work, we formulate the SFT and alignment processes for LLMs as one of constrained optimization: we minimize the task perplexity while simultaneously imposing custom constraints on preferences. This enables customization of the LLM to different preferential properties while maintaining performance on the original task. Consequently, we propose Lagrange Large Language Models (L3Ms) to solve this constrained problem by incorporating the constraints in the objective function using the logarithmic barrier. We include experimental results to illustrate the customization qualities of L3Ms, which can fit to different preferences, providing a personalized user experience.
@@ Appendix A Experimental setup
In addition to Section 6.1, here we provide further details of the experimental setup.
@@
We use UltraChat (Ding et al., 2023), a large-scale dataset of instructional conversations, as our task data to induce instruction-following capabilities. Since each sample contains a sequence of multi-turn question-answer pairs, we randomly sample one of the answers as the response and treat the preceding dialogue as the prompt. We then filter such (prompt, response) pairs to a maximum token length of 512. Consequently, we obtain 340k training samples, 1.7k validation samples, and 1.7 test samples, split randomly as the dataset does not contain train-val-test splits.
@@
We train LLMs for 1 epoch on the task data, with a mini-batch size of 64. We use Adam with a learning rate of 10-6, and a cosine learning rate scheduler (with 5% of the epoch used for warmup). We set weight decay to 0.1 and the gradient clipping maximum norm to 1. We also utilize 16-bit (mixed) precision training and gradient checkpointing. We exponentially decay the log-barrier parameter $\mu$ during training from 1 to 10-6, and use a smoothing factor of 0.1 for the exponential moving average. Lastly, we use top-$p$ sampling ($p$ set to 0.9) for response generation. Apart from this, we use the default hyper-parameters in the Transformers library (Wolf et al., 2020).
@@ A.1 Learning preference reward models
While some preference reward functions are engineered or rule-based, others are learned. Such preferences can often be difficult to quantify. Alternatively, it is easier to compare responses with respect to the preference, e.g., ranking them from most to least helpful. Consequently, the data for learning preference reward models consist of tuples of the form $(x,y_{+},y_{-})$, where the prompt $x$ is accompanied by two responses $y_{+}$ and $y_{-}$, with the former response being preferred over the latter.
@@
Taking inspiration from Rafailov et al. (2023), we initialize the preference reward model $r_{\phi}(\cdot)$ as a pre-trained LLM, and set the reward to be its length-normalized log-likelihood. In this way, we utilize the pre-trained model fully, not just its backbone. As the preference reward model is fine-tuned, its log-likelihoods are updated to differentiate the preferred responses from the rejected ones.
@@
We use the Helpful and Harmless (Bai et al., 2022) preference data to learn two reward models respectively. We obtain 161k training samples and 9k test samples after filtering the (prompt, response) pairs to a maximum token length of 2024; 3/4-th are for helpfulness while the remaining 1/4-th are for harmlessness. We further set 5% of the training data for validation.
@@
We initialize all preference reward functions with LLaMA-7B. We train for 2 epochs with a mini-batch size of 64. We use Adam with a learning rate of 10-6, and a cosine learning rate scheduler (with 10% of the epoch used for warmup). We set weight decay to 0.1 and the gradient clipping maximum norm to 1. We also utilize 16-bit (mixed) precision training and gradient checkpointing. Apart from this, we use the default hyper-parameters in the Transformers library (Wolf et al., 2020). We validate after every 10% of the epoch and save the best model.

### v2

@@ 5 Lagrange Large Language Models (L3Ms)
Thus far, we have formulated the SFT and alignment stages as a constrained optimization problem in Eq. 5. We proceed to find solutions for the same by solving the unconstrained objective in Eq. 10. We call the family of models obtained in this way L3Ms, i.e., Lagrange Large Language Models.
@@ 5.1 Optimization procedure
Since the log barrier converges to the characteristic function as $\mu\rightarrow 0$, we want to find the minimizer of $\mathcal{G}_{\mu}(\theta)$ for a very small $\mu$. However, doing so directly leads to instabilities as the objective function is ill-conditioned. Instead, it is common practice to follow an iterative procedure: one finds the minimizer for a fixed $\mu$, reduces $\mu$, and repeats (Curtis et al., 2024). Specifically, the procedure is instantiated with initial values $\theta_{0}$, $\mu_{0}$, and $10<\gamma<$ 1. On the $t$-th iteration, $\mu_{t}\leftarrow\gamma\mu_{t-1}$ is reduced and $\theta_{t}\leftarrow\arg_{\theta}\min\mathcal{G}_{\mu_{t}}(\theta)$ (with initialization $\theta_{t-1}$). In doing so, the constraints are gradually enforced, nudging the LLM to satisfy them over the optimization procedure while avoiding instabilities. As $\{\mu_{t}\}\searrow 0$, the weights $\{\theta_{t}\}$ converge to the minimizer of the constrained problem.
It is impossible to minimize $\mathcal{G}_{\mu_{t}}(\theta)$ exactly in many practical applications. Instead, at each iteration, one can take a single optimization step toward the solution. Doing so is amenable to stochastic gradient methods and mitigates computational overhead: the optimization proceeds as normal while the value of $\mu$ is reduced over the course of the procedure. One can guarantee the convergence of this procedure to the optimal solution in some settings; for example, Curtis et al. (2024) prove convergence when dealing with box constraints. However, convergence in a scenario like ours is not guaranteed. Nevertheless, we will experimentally demonstrate its use for our constrained problems.
@@
This follows immediately from Eqs. 9 and 10. Note that the gradients $\partial_{\theta}C_{i}(\theta)$’s are also known as policy gradients in reinforcement learning literature. We discuss our strategy for estimating these gradients in Appendix B and refer readers to Schulman et al. (2016) for a more detailed review.
@@ 5.2 Connection to Lagrange multipliers
The log barrier and the Lagrangian are intrinsically connected; this becomes evident when comparing Eq. 11 with the (gradient of the) Lagrangian in Eq. 8. In particular, we define the multipliers:
@@
5.3 Memory and time complexity
L3Ms differ from traditional LLMs only in the fine-tuning process. In fact, L3Ms require less memory and are faster to fine-tune compared to traditional LLMs. This comes from merging the SFT and alignment stages as we did in our constrained optimization formulation in Eq. 5. We minimize the task objective directly and avoid the need to compute deviations away from the SFT model (for instance, to impose the KL divergence penalty). As a result, we save on loading the SFT model into memory and evaluating it on generated responses at each optimization step. Also, the L3M’s log barrier parameter $\mu$ is adjusted during fine-tuning itself, without needing any extra steps.
@@ 5.4 Implementation details
Alternating the objective and gradient clipping
Gradient clipping is a simple yet effective way to ensure stable training of large models (Goodfellow et al., 2016; Zhang et al., 2020). We employ this technique, albeit with the modification of clipping the gradients of both the task objective and the constraints separately, as they can have varying magnitudes. We achieve this by alternating between reducing the task objective and enforcing the constraints by flipping a fair coin to select one or the other. While this doubles the number of steps to achieve the same effect, it does not increase the amount of work done as now only one part of the objective or the other is evaluated at each step.
Length normalization
The gradient of our objective function in Eq. 11 involves the LLM’s log-likelihoods on the generated responses through the gradients $\partial_{\theta}C_{i}(\theta)$’s (cf. Eq. 12). To avoid a response length bias, we length-normalize the log-likelihoods, akin to the definition of perplexity.
Estimating the mean preference rewards
We need to estimate the expectations involved in the gradient of our objective function in Eq. 11. The expectations in the numerators can be estimated with the per-mini-batch Monte Carlo averages (Mohamed et al., 2020). However, $C_{i}(\theta)$’s in the denominator need careful consideration. Note that: (i) $C_{i}(\theta)$ does not involve the gradient, so its estimate can include information from previous mini-batches to reduce the estimate’s variance, and (ii) since the weight $\theta$ is updated during fine-tuning, $C_{i}(\theta)$ is non-stationary. Hence, we use an exponential moving average estimate for the mean (offset) preference rewards $C_{i}(\theta)$ in the denominator.
@@ 6 Experimental results
To illustrate the customization of L3Ms, we empirically evaluate them on: (i) satisfaction of the imposed constraints, and (ii) minimization of the task objective. Our code, based on the Transformers library (Wolf et al., 2020), is available at: https://github.com/Guneet-Dhillon/l3m.
@@ 6.1 Setup
We use LLaMA-7B (Touvron et al., 2023) for all our experiments, as it is a lightweight LLM pre-trained on a large corpus. We are interested in the task of instruction-following, for which we use UltraChat (Ding et al., 2023), a large-scale dataset of instructional conversations. We run all experiments on NVIDIA H100s. Further details of our experimental setup are included in Appendix A.
We refer to the LLM fine-tuned to minimize the SFT objective only (without the alignment stage) as the SFT model. We fine-tune LLMs using the minimax approach to find a saddle point of the Lagrangian, as proposed by Moskovitz et al. (2024); Dai et al. (2024), and refer to them as MMs. Lastly, we refer to the LLMs fine-tuned using our proposed approach as L3Ms.
@@ 6.2 Length constrained L3Ms
Consider tasks in which the lengths of the responses need to be contained in the range $[l_{\text{low}},l_{\text{high}}]$ to control verbosity; for example, in summarization tasks (Makino et al., 2019). In this case, the natural choice for the reward functions compute the response length and its negation: $r_{1}(y|x)=|y|$ and $r_{2}(y|x)=-|y|$. Furthermore, these rewards are to be controlled with the minimum requirements of $l_{\text{low}}$ and $-l_{\text{high}}$, respectively. Note that these reward functions are perfectly negatively correlated.
If we naively average the rewards, any unconstrained formulation of alignment (including RLHF) will be ineffective, as the loss will always vanish due to the anti-correlation. We could use a weighted average and tune the weights heuristically, but this is tedious. Instead, we use the constrained formulation and directly constrain the rewards $r_{1}(y|x)=|y|\geq l_{\text{low}}$ and $r_{2}(y|x)=-|y|\geq-l_{\text{high}}$.
We fine-tune several L3Ms with varying length constraints. We illustrate the distributions of the generated response lengths (in tokens) and report the perplexities achieved on the task-related data in Fig. 2. We observe that the mean response lengths are in the required range in each case, satisfying the imposed average constraints. Additionally, the task perplexities increase slightly as the constraints on the response lengths are made more stringent. However, there is little to no degradation relative to the SFT model, with all mean task perplexities being within 0.02 standard deviations.
The examples included in Table 1 are generated from such L3Ms. While all the responses correctly answer the prompts, their lengths vary, corresponding to the length constraints imposed on them.
@@
Figure 2: Length constrained L3Ms. We report the response lengths (in tokens) and task perplexities of the SFT model and the L3Ms with varying length constraints. Left: The mean response length with the mean and standard deviation of the task perplexities. Right: The distribution of the response lengths. The notches indicate the medians and their 95% confidence intervals, the boxes show the $\pm$25quantiles, and the whiskers denote the 1.5$\times$ interquartile ranges. The white circles mark the means, and the black dashed lines depict the constraints imposed on the different L3Ms.
@@ 6.3 Helpful and harmless L3Ms
Next, we consider the Helpful and Harmless (HH) preferences that have been extensively used in the LLM alignment literature (Ji et al., 2023; Wang et al., 2024; Zhou et al., 2024; Guo et al., 2024). Specifically, we utilize the datasets by Bai et al. (2022) to train two preference reward functions, respectively. These learned reward functions are negatively correlated (Bai et al., 2022; Dai et al., 2024). Furthermore, note that the numerical outputs of both these reward functions are interpreted as ratings such that a higher numerical value indicates higher helpfulness/harmlessness and vice versa.
We fine-tune several L3Ms with varying HH constraints. We compare our L3M approach of using log barriers with that of the minimax optimization by Moskovitz et al. (2024); Dai et al. (2024) to find a saddle point of the Lagrangian (MMs). In our experience, learning the Lagrange multipliers with the latter is extremely sensitive to the choice of the learning rate. Moreover, to avoid loading an additional SFT model during fine-tuning, as is done by Moskovitz et al. (2024); Dai et al. (2024), our implementation of MM minimizes the task objective directly (as is done by L3Ms as well).
Fig. 3 shows the achieved task perplexities and helpful-harmless rewards for the different LLMs. At initialization, the helpful-harmless rewards are both low, with a high task perplexity of 1.316$\pm$0.4. This is improved upon by the SFT model, reducing the perplexity to 0.805$\pm$0.3 and attaining a high helpfulness reward (due to the task data instilling instruction-following capabilities). Furthermore, MMs sacrifice task performance to oversatisfy the constraints, with mean task perplexities $\geq$ 0.822. Conversely, L3Ms satisfy the imposed helpful-harmless reward constraints with consistently lower task perplexities (the mean perplexities are in the range 0.804-0.820). We attribute this to the L3Ms having better Lagrange multipliers by design rather than learning them as in MMs (cf. Section 5.2).
@@ 7 Conclusions
In this work, we formulate SFT and alignment in LLMs as constrained optimization: we minimize the task objective while simultaneously imposing application-specific constraints on preferences. This enables the customization of LLMs to different preferential properties while maintaining performance on the task of interest. Consequently, we propose Lagrange Large Language Models (L3Ms) to solve this constrained optimization problem by incorporating the constraints in the objective using the logarithmic barrier. We include experimental results to illustrate the customization qualities of L3Ms, which can fit to different preferences, providing a personalized user experience.
@@ Appendix A Experimental setup
In addition to the experimental setup discussed in Section 6.1, we provide further details here.
@@
We use UltraChat (Ding et al., 2023), a large-scale dataset of instructional conversations, as our task data to induce instruction-following capabilities. Since each sample contains a sequence of multi-turn question-answer pairs, we randomly sample one of the answers as the response and treat the preceding dialogue as the prompt. We then filter such (prompt, response) pairs to a maximum token length of 512. Consequently, we obtain 340k training samples, 1.7k validation samples, and 1.7 test samples, split randomly since the dataset does not contain train-val-test splits.
@@
We fine-tune LLMs for 1 epoch on the task data, with a mini-batch size of 64. We use Adam with a learning rate of 10-6 and a cosine learning rate scheduler (with 5% of the epoch used for warmup). We set weight decay to 0.1 and the gradient clipping maximum norm to 1. We utilize 16-bit (mixed) precision training and gradient checkpointing. We exponentially decay the log-barrier parameter $\mu$ during fine-tuning from 1 to 10-6 and use a smoothing factor of 0.1 for the exponential moving average. Lastly, we use top-$p$ sampling ($p$ set to 0.9) for response generation. Apart from this, we use the default hyper-parameters in the Transformers library (Wolf et al., 2020).
@@ A.1 Learning preference reward models
While some preference reward functions are engineered or rule-based, others are learned. Such preferences can often be difficult to quantify. Alternatively, it is easier to compare responses with respect to the preference, e.g., ranking them from most to least helpful. Consequently, the data for learning preference reward models consist of tuples of the form $(x,y_{+},y_{-})$, where the prompt $x$ is accompanied by two responses $y_{+}$ and $y_{-}$, with a preference for the former response over the latter.
@@
Taking inspiration from Rafailov et al. (2023), we initialize the preference reward model $r_{\phi}(\cdot)$ as a pre-trained LLM and set the reward to be its length-normalized log-likelihood. In this way, we utilize the pre-trained model fully, not just its backbone. As the preference reward model is fine-tuned, its log-likelihoods/rewards are updated to differentiate the preferred responses from the rejected ones.
@@
We use the Helpful and Harmless (Bai et al., 2022) preference data to learn two reward models, respectively. We obtain 161k training samples and 9k test samples after filtering the (prompt, response) pairs to a maximum token length of 2024; 3/4-th are for helpfulness while the remaining 1/4-th are for harmlessness. We further set 5% of the training data for validation.
@@
We initialize all reward models with LLaMA-7B (Touvron et al., 2023). We fine-tune for 2 epochs with a mini-batch size of 64. We use Adam with a learning rate of 10-6 and a cosine learning rate scheduler (with 10% of the epoch used for warmup). We set weight decay to 0.1 and the gradient clipping maximum norm to 1. We utilize 16-bit (mixed) precision training and gradient checkpointing. Apart from this, we use the default hyper-parameters in the Transformers library (Wolf et al., 2020). We validate after every 10% of the epoch and save the best model.